question : find general solution for given equation: sin 3x + cos 2x = 0 .

question     :     find general solution for given equation:    sin 3x  +  cos 2x  =  0                .

Grade:12

1 Answers

Arun
6 years ago
Dear Ankur
 
cos 2x = - sin 3x
 
cos 2x = cos (\pi/2 + 3x)
 
Now
 
2x = 2n\pi ± (\pi/2 + 3x)
 
Taking +ve sign
 
2x = 2n\pi + \pi/2 + 3x
 
x = -(2n\pi + \pi/2)
 
 
Now taking -ve sign
 
2x = 2n\pi - \pi/2 - 3x
 
x = (1/5)(2n\pi - \pi/2)
 
 
 
Regards
Arun (askIITians forum expert)